# Probabilistic Approaches to The Energy Equality in Forced Surface Quasi-Geostrophic Equations

### v1

@@ 1.2. Key idea and technical comment
Large deviations. Concerning the large deviations of (1.3), the main obstacles arise from two aspects. On the one hand, generalized Leray solutions of (1.3) are probabilistically weak and therefore cannot be represented as measurable maps of Brownian paths. This prevents us from using the well-known weak convergence approach by Budhiraja, Dupuis, and Maroulas [BDM08] for large deviations of stochastic PDEs. On the other hand, the uniqueness of the skeleton equation (1.1) is unknown in such critical and supercritical ($\alpha\in(0,1/2]$) regimes. Large deviations for subcritical stochastic SQG equations driven by multiplicative noise have been studied by Liu, Röckner and Zhu in [LRZ13], under the assumption that $\theta_{0}\in L^{p}\cap H^{\delta}$, where $\delta>2-2\alpha$ and $1/p\in(0,\alpha-1/2)$. In this case, the pathwise uniqueness of the stochastic SQG equations is obtained, and the skeleton equation also has a unique solution, which leads to the weak convergence approach. In general, the well-posedness of the skeleton equation plays a key role in large deviations. We refer readers to [Hey23] for more details, where a counterexample is proposed to illustrate the violation of the lower bound due to the lack of uniqueness. Also see [FG23, Section 8] for the consistency of the rate function and its lower semi-continuous envelope with the help of the uniqueness and stability of the skeleton equation. These two obstacles have also been encountered in the study of three-dimensional Landau-Lifshitz-Navier-Stokes equations and were resolved by Gess, Heydecker, and the second author [GHW24]. To summarize the key idea, the upper and lower bounds are investigated separately, with the lower bound being obtained manually by restricting to the weak-strong uniqueness class. In this context, we will adopt the same idea as in [GHW24] on the probabilistic side, but extend it on the analytic side. Specifically, we generalize this argument into an $\dot{H}^{-1/2}$-framework.
@@
where $H_{j}\theta$ is the high-frequency part of $\theta$ and $S_{j}\theta$ is the low-frequency part defined by (2.1) in Section 2. It has been shown by Marchand in [Mar08a, Lemma 9.3] that the first two terms on the right-hand side of (1.9) are easier to estimate due to the regularizing effect of $S_{j}$. However, the estimation of $\nabla\cdot\left(H_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right)$ is more challenging and requires the commutator estimates. In particular, for every $\phi\in C^{\infty}(\mathbb{T}^{2})$, $\iota>0$ and $r\in(-1,1/2],$ we have
(${}_{\dot{H}^{-3-r-\iota}}\left\langle\nabla\cdot\left(H_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right),\phi\right\rangle_{\dot{H}^{3+r+\iota}}\leqslant C\|\phi\|_{\dot{H}^{3+r+\iota}(\mathbb{T}^{2})}\left\|H_{j}\theta\right\|_{\dot{H}^{-r}(\mathbb{T}^{2})}^{2}.$
The energy equality. In this part, we summarize the key idea of the proof for Theorem 1.4. The main idea arises from [GH23, Section 11] and [GHW24, Section 9]. In particular, in [GHW24, Section 9], the authors provide a purely analytic approach concerning the relationship between the energy equality and the weak-strong uniqueness class for three-dimensional forced Navier-Stokes equations. Moreover, they provide a brief explanation of its probabilistic interpretation. In this paper, we adopt the probabilistic approach for proving this relationship when $\beta=\alpha/2$.
@@
both $\theta_{\varepsilon,m}$ and its time-reversal process $\mathfrak{T}_{T}\theta_{\varepsilon,m}$ satisfy large deviations as presented in Theorem 1.3. On the domain where the large deviations lower bound matches the upper bound ($\mathcal{C}$ in Theorem 1.3), the uniqueness of the large deviations rate function holds. Therefore, we have that $\mathcal{I}(\mathfrak{T}_{T}\theta)=\mathcal{I}(\theta)$ for all $\theta\in\mathbb{X}_{\alpha,\alpha/2}$ satisfying $\theta,\mathfrak{T}_{T}\theta\in\mathcal{C}$. For any weak solution $\theta$ of (1.1), direct computation shows that
$\mathcal{I}(\theta)=\|\theta(0)\|_{L^{2}(\mathbb{T}^{2})}^{2},\ \ \mathcal{I}(\mathfrak{T}_{T}\theta)=\|\theta(T)\|_{L^{2}(\mathbb{T}^{2})}^{2}+2\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s-2\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s.$
Then the kinetic energy equality holds as long as $\theta,\mathfrak{T}_{T}\theta\in\mathcal{C}$. Based on this observation, one can see that the uniqueness of the rate function plays a key role. Since the uniqueness holds on the domain where the lower bound matches the upper bound, then one can provide a sufficient condition concerning the open problem of the validity of the kinetic energy equality for arbitrary Leray solutions. As long as one can prove a full large deviation of (1.10) on $\mathbb{X}_{\alpha,\alpha/2}$ with respect to the rate function $\mathcal{I}$, then this implies that the kinetic energy equality holds on $\mathbb{X}_{\alpha,\alpha/2}$ without restriction.
@@ 1.3. Comments of literature
Large deviations for stochastic PDEs in singular limits. Large deviations of stochastic scalar conservation laws in the joint limits of vanishing noise and viscosity were studied by Mariani [Mar10]. In the work of Dirr, Fehrman and Gess [DFG24], large deviations for a conservative stochastic PDE corresponding to the simple symmetric exclusion process were established. Ferman and Gess [FG23] demonstrated that a class of generalized Dean-Kawasaki type equations satisfies large deviations with the same rate function as that of the zero range process. The second author and Zhang [WZ24] extended the analysis to a more general case concerning singular interactions for the Dean-Kawasaki equation. Gess, Heydecker, and the second author [GHW24] proved the large deviation principle for the Landau-Lifshitz-Navier-Stokes equations, whose rate function is consistent with the Quastel-Yau lattice gas model [QY98]. Further works concerning large deviations for stochastic PDEs in scaling limits with vanishing noise intensity and correlations can be found in [HW15, CD19a, CD19b, CP22].
Background of the energy functions. There are two types of energy that have physical significance for the SQG equation: the Hamiltonian and the kinetic energy. For the inviscid SQG equation, the functional
can be taken as the Hamiltonian of the system. This is derived from the Euler-Poincaré variational principle. We refer readers to Arnold’s work [Arn66], which is an infinite-dimensional generalization of [Poi01]. (See [Res95, Section 2.2], [BSV19, Section 1.4, Appendix A.1] for more details.) Besides the Hamiltonian $\mathcal{H}$, the kinetic energy is widely discussed as well. In the system of the SQG equation, the field $\theta$ represents the temperature or surface buoyancy for a rapidly rotating stratified fluid with uniform potential vorticity, and $u_{\theta}=R^{\perp}\theta$ denotes the transport velocity field [HPGS95]. Since the Riesz transform is an $L^{2}$-isometry [SW71, Chapter VI, Section 2], the kinetic energy can be represented by
The validity of the energy equality of the SQG equation with respect to both of these two kinds of energy has been widely investigated.
@@
Energy equality for the SQG equation. Energy conservation of the inviscid SQG equation is related to Onsager’s conjecture [Ons49]. Isett and Vicol proved in [IV15] that $\theta\in L^{3}_{t}L^{3}_{x}$ implies the conservation of the Hamiltonian $\mathcal{H}$. For sufficiently regular solutions, the $L^{2}$-norm (the kinetic energy) is conserved as well. Several Besov-type regularity conditions are proposed for weak solutions to guarantee the conservation of the kinetic energy, see [Zho05], [Cha06a], [WYY23]. Dai [Dai17] investigated the kinetic energy equality for viscosity solutions of the super-critical dissipative SQG equation under some regularity conditions.
@@
Weak-strong uniqueness for the SQG equation. Results of weak-strong uniqueness for the SQG equations with different fractional dissipation index $\alpha$ and different initial data $\theta_{0}$ can be found in the following literature. Constantin and Wu [CW99] proved the uniqueness of strong solution when $\alpha\in(1/2,1]$ with $\theta_{0}$ in $\dot{H}^{-1/2}$ or $L^{2}$ , which is analogous to the Ladyzhenskaya-Prodi-Serrin condition of Navier-Stokes equation [KL57, Pro59, Ser62]. In the critical and supercritical cases $\alpha\in(0,1/2]$, there are several results for $L^{2}$-valued initial data. Dong and Chen in [DC06] proposed Ladyzhenskaya-Prodi-Serrin type condition $L^{r}_{t}L^{p}_{x}$ to $\nabla\theta$ and showed a weak-strong uniqueness property. In [DC12], the authors extended this regularity condition to Besov space $L^{r}_{t}B^{0}_{p,\infty}$. For the critical case $\alpha=1/2$, Marchand [Mar08b] proved the weak-strong uniqueness in BMO-type space. We will summarize examples of weak-strong uniqueness for subcritical, critical, and supercritical cases in Appendix A, and illustrate that these regularity classes can be taken as examples of the domain where we restrict the lower bound of large deviations. Furthermore, we point out that, to the best of our knowledge, weak-strong uniqueness results for the supercritical case $\alpha\in(0,1/2)$ with $\theta_{0}$ in $\dot{H}^{-1/2}$ have not been obtained. In this case, the dissipation term is not sufficient to prevent weak solutions from being distribution-valued. Therefore, it is challenging to propose a reasonable regularity condition for proving the uniqueness.
@@
Relationship to the study of convex integration. From the previous explanations, it can be inferred that obtaining matching upper and lower bounds for large deviations is related to the uniqueness of solutions for (1.1). Recently, the convex integration technique has led to various non-uniqueness results for fluid equations. Buckmaster, Shkoller, and Vicol [BSV19] demonstrated the non-uniqueness of weak solutions for both inviscid and dissipative SQG equations by converting the equation into momentum form. Subsequently, different convex integration schemes have been employed to construct non-unique solutions for inviscid and dissipative SQG equations, as illustrated in [CKL21] and [IM21]. All of these works focus on the SQG equations without external forces, where the solutions satisfy $\Lambda^{-1}\theta\in C_{t}^{\sigma}C_{x}^{s}$ for some $\sigma$ and $s$. However, it remains an open question whether $\theta\in L^{2}_{t}H^{\alpha}_{x}$, which is the space considered in Theorem 1.4. For example, in [BSV19], the solutions were constructed in $C_{t}^{\sigma}C_{x}^{s}$ with $51/2<s<4/$ 5 and $\sigma<s/(2-s).$ Therefore, these non-uniqueness results do not imply a violation of the full large deviations.
@@
In the context of stochastic SQG equations, Yamazaki established the non-uniqueness of solutions driven by trace-class noise [Yam22], and later extended these results to include linear multiplicative noise [Yam23]. In [HZZ23] and [HLZZ24], the authors discussed the cases of irregular spatial noise and space-time white noise, respectively. We remark that, in [HLZZ24], infinitely many non-Gaussian ergodic stationary solutions of (1.3) (with $\delta=0$) in $B^{-1/2}_{p,1},p\geqslant 2,$ were constructed. We hope to connect the discussion of non-Gaussian stationary solutions with the uniqueness problem of deterministic SQG equations through the non-Gaussian large deviations in Section 8.
@@ 1.4. Structure of the paper
This paper is organized as follows. In Section 2, we introduce some preliminaries which will be used throughout the paper. Section 3 is dedicated to proving the existence and exponential tightness for stochastic generalized Leray solutions. Section 4 presents a variational characterization of the rate function. In Section 5, we establish the upper bound for large deviations. Section 6 provides a restricted lower bound for large deviations using the entropy method. In Section 7, we prove a relationship between the weak-strong uniqueness class and the energy equality for (1.1). Finally, in Section 8, we discuss an explicit representation of the quasi-potential. Additionally, Appendix A summarizes several examples of weak-strong uniqueness classes concerning different ranges of parameters $\alpha$ and $\beta$, and Appendix B demonstrates the large deviations for Gaussian initial conditions.
@@ 2.1. Notations and spaces
endowed with the norm
$\|f\|_{W^{\sigma,p}([0,T];E)}^{p}:=\int_{0}^{T}\|f(t)\|_{E}^{p}\mathrm{\,d}t+\int_{0}^{T}\int_{0}^{T}\frac{\|f(t)-f(s)\|_{E}^{p}}{|t-s|^{1+\sigma p}}\mathrm{\,d}t\,\mathrm{d}s.$
We denote the distribution space on $\mathbb{T}^{2}$ by $\mathscr{D}^{\prime}(\mathbb{T}^{2})$, which is the dual space of $C^{\infty}(\mathbb{T}^{2})$. Let $f\in\mathscr{D}^{\prime}(\mathbb{T}^{2})$ and let $g=(g(k))_{k\in\mathbb{Z}^{2}}$ be at most of polynomial growth. The Fourier transform of $f$ and the inverse Fourier transform of $g$ are defined as
$\mathcal{F}_{\mathbb{T}^{2}}f(k):=\hat{f}(k):=\int_{\mathbb{T}^{2}}e^{2\pi\mathrm{i}x\cdot k}f(x)\mathrm{\,d}x,\quad k\in\mathbb{Z}^{2},\quad\mathcal{F}_{\mathbb{T}^{2}}^{-1}g(x):=\sum_{k\in\mathbb{Z}^{2}}e^{2\pi\mathrm{i}x\cdot k}g(k),\quad x\in\mathbb{T}^{2}.$
@@
is an orthonormal eigenbasis of $\Lambda$ on $H$, which we denote by $\{e_{k}\}$. For every $s>0$, define
$\|f\|_{H^{s}(\mathbb{T}^{2})}^{2}:=\sum_{k\in\mathbb{Z}^{2}}|2\pi k|^{2s}\big{|}\hat{f}(k)\big{|}^{2},$
and let $H^{s}(\mathbb{T}^{2})$ denote the space of all $f\in H$ for which $\|f\|_{H^{s}(\mathbb{T}^{2})}$ is finite. Define $H^{-s}(\mathbb{T}^{2})$ to be the dual of $H^{s}(\mathbb{T}^{2})$. Finally, let $H^{\infty}(\mathbb{T}^{2}):=\bigcap\limits_{s>0}H^{s}(\mathbb{T}^{2})$.
Now we recall the Littlewood-Paley decomposition briefly. Let $\varphi\in C_{c}^{\infty}\left(\mathbb{R}^{2}\right)$ be a non-negative radial function so that $\varphi(\xi)=1$ for $|\xi|\leqslant 1/2$ and $\varphi(\xi)=0$ for $|\xi|\geqslant 1$. Let $\psi(\xi)=\varphi(\xi/2)-\varphi(\xi)$. Then $\big{(}\psi(\cdot/2^{j})\big{)}_{j\in\mathbb{Z}}$ is a dyadic partition of unity. For every $j\in\mathbb{Z}$, we define the $j$-th dyadic block of the Littlewood-Paley decomposition of $f\in\mathscr{D}^{\prime}(\mathbb{T}^{2})$ by
$\Delta_{j}f:=\mathcal{F}_{\mathbb{T}^{2}}^{-1}\left(\psi\left(\cdot/2^{j}\right)\hat{f}\right).$
The low-frequencies and high-frequencies cutting-off operators are defined by
$S_{j}f:=\mathcal{F}_{\mathbb{T}^{2}}^{-1}\left(\varphi\left(\cdot/2^{j}\right)\hat{f}\right)=\sum_{k=-\infty}^{j-1}\Delta_{k}f,$ (2.1)
$H_{j}f:=\left(Id-S_{j}\right)f=\sum_{k=j}^{\infty}\Delta_{k}f.$
For every $\phi\in H^{\infty}(\mathbb{T}^{2})$, the commutator between $\Lambda$ and $\phi$ is defined by $[\Lambda,\phi]g:=\Lambda(\phi g)-\phi\Lambda g$, for any $g\in H^{\infty}(\mathbb{T}^{2}).$
The following lemma enables us to introduce the commutator estimate, which will contribute to the well-definedness of the nonlinear term.
@@
Lemma 2.1 can be proved by integration by parts, see for example, [Mar08a, Lemma 2.1].
@@
This commutator estimate is stated for $\mathbb{R}^{2}$ functions in [Mar08a, Lemma 2.2]. The same proof also works in the case of $\mathbb{T}^{2}$ functions, see [ZZ14, Lemma 4.2], [BSV19, Lemma A.5].
@@
$\left\langle\nabla\cdot\left(H_{j}fR^{\perp}\left(H_{j}f\right)\right),\phi\right\rangle:=$ $-{}_{H^{-1/2}}\big{\langle}\frac{1}{2}R_{2}\left(H_{j}f\right),\left[\Lambda,\partial_{1}\phi\right]\Lambda^{-1}\left(H_{j}f\right)\big{\rangle}_{H^{1/2}}$ (2.3)

### v3

@@ 1.2. Key idea and technical comment
Large deviations. Concerning the large deviations of (1.3), the main obstacles arise from two aspects. On the one hand, generalized Leray solutions of (1.3) are probabilistically weak and therefore cannot be represented as measurable maps of Brownian paths. This prevents us from using the well-known weak convergence approach by Budhiraja, Dupuis, and Maroulas [BDM08_LDP] for large deviations of stochastic PDEs. On the other hand, the uniqueness of the skeleton equation (1.1) is unknown in such critical and supercritical ($\alpha\in(0,1/2]$) regimes. Large deviations for subcritical stochastic SQG equations driven by multiplicative noise have been studied by Liu, Röckner, and Zhu in [LRZ13_LDP], under the assumption that $\theta_{0}\in L^{p}\cap H^{\delta}$, where $\delta>2-2\alpha$ and $1/p\in(0,\alpha-1/2)$. In this case, the pathwise uniqueness of the stochastic SQG equations is obtained, and the skeleton equation also has a unique solution, which leads to the weak convergence approach. In general, the well-posedness of the skeleton equation plays a key role in large deviations. We refer readers to [Hey23] for more details, where a counterexample is proposed to illustrate the violation of the lower bound due to the lack of uniqueness. Also see [FG23, Section 8] for the consistency of the rate function and its lower semi-continuous envelope with the help of the uniqueness and stability of the skeleton equation. These two obstacles have also been encountered in the study of three-dimensional Landau-Lifshitz-Navier-Stokes equations and were resolved by Gess, Heydecker, and the second author [GHW24_LLNS]. To summarize the key idea, the upper and lower bounds are investigated separately, with the lower bound being obtained manually by restricting to the weak-strong uniqueness class. In this context, we will adopt the same idea as in [GHW24_LLNS] on the probabilistic side, but extend it on the analytic side. Specifically, we generalize this argument into an $\dot{H}^{-1/2}$-framework.
@@
where $H_{j}\theta$ is the high-frequencies part of $\theta$ and $S_{j}\theta$ is the low-frequencies part defined by (2.1) in Section 2. It has been shown by Marchand in [Mar08_existence, Lemma 9.3] that the first two terms on the right-hand side of (1.11) are easier to estimate due to the regularizing effect of $S_{j}$. However, the estimation of $\nabla\cdot\left(H_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right)$ is more challenging and requires the commutator estimates. In particular, for every $\phi\in C^{\infty}(\mathbb{T}^{2})$, $\iota>0$ and $r\in(-1,1/2],$ we have
$_{\dot{H}^{-3-r-\iota}}\left\langle\nabla\cdot\left(H_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right),\phi\right\rangle_{\dot{H}^{3+r+\iota}}\leqslant C\|\phi\|_{\dot{H}^{3+r+\iota}(\mathbb{T}^{2})}\left\|H_{j}\theta\right\|_{\dot{H}^{-r}(\mathbb{T}^{2})}^{2}.$
The energy equality. In this part, we summarize the key idea of the proof for Theorem 1.4. The main idea arises from [GH23, Section 11] and [GHW24_LLNS, Section 9]. In particular, in [GHW24_LLNS, Section 9], the authors provide a purely analytic approach concerning the relationship between the energy equality and the weak-strong uniqueness class for three-dimensional forced Navier-Stokes equations. Moreover, they provide a brief explanation of its probabilistic interpretation. In this paper, we adopt the probabilistic approach for proving this relationship when $\beta=\alpha/2$.
@@
both $\theta_{\varepsilon,m(\varepsilon)}$ and its time-reversal process $\mathfrak{T}_{T}\theta_{\varepsilon,m(\varepsilon)}$ satisfy large deviations as presented in Theorem 1.3. On the domain where the large deviations lower bound matches the upper bound ($\mathcal{C}$ in Theorem 1.3), the uniqueness of the large deviations rate function holds. Therefore, we have that $\mathcal{I}(\mathfrak{T}_{T}\theta)=\mathcal{I}(\theta)$ for all $\theta\in\mathbb{X}_{\alpha,\alpha/2}$ satisfying $\theta,\mathfrak{T}_{T}\theta\in\mathcal{C}$. For any weak solution $\theta$ of (1.1), direct computation shows that
$\mathcal{I}(\theta)$ $=\|\theta(0)\|_{L^{2}(\mathbb{T}^{2})}^{2},$
$\mathcal{I}(\mathfrak{T}_{T}\theta)$ $=\|\theta(T)\|_{L^{2}(\mathbb{T}^{2})}^{2}+2\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s-2\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s.$
Then the kinetic energy equality holds as long as $\theta,\mathfrak{T}_{T}\theta\in\mathcal{C}$. Based on this observation, one can see that the uniqueness of the rate function plays a key role. Since the uniqueness holds on the domain where the lower bound matches the upper bound, then one can provide a sufficient condition concerning the open problem of the validity of the kinetic energy equality for arbitrary Leray solutions. As long as a full large deviation of (1.12) on $\mathbb{X}_{\alpha,\alpha/2}$ with respect to the rate function $\mathcal{I}$ was proved, then this implies that the kinetic energy equality holds on $\mathbb{X}_{\alpha,\alpha/2}$ without restriction.
@@ 1.3. Comments of literature
Large deviations for stochastic PDEs in singular limits. Large deviations of stochastic scalar conservation laws in the joint limits of vanishing noise and viscosity were studied by Mariani [Mariani10]. In the work of Dirr, Fehrman, and Gess [DFG24], large deviations for a conservative stochastic PDE corresponding to the simple symmetric exclusion process were established. Fehrman and Gess [FG23] demonstrated that a class of generalized Dean-Kawasaki type equations satisfies large deviations with the same rate function as that of the zero range process. The second author and Zhang [WZ24] extended the analysis to a more general case concerning singular interactions for the Dean-Kawasaki equation. Gess, Heydecker, and the second author [GHW24_LLNS] proved the large deviation principle for the Landau-Lifshitz-Navier-Stokes equations, whose rate function is consistent with the Quastel-Yau lattice gas model [QY98]. Further works concerning large deviations for stochastic PDEs in scaling limits with vanishing noise intensity and correlations can be found in [HW15_LDP, CD19_LDP1, CD19_LDP2, CP22].
@@
Energy equality for the SQG equation. Energy conservation of the inviscid SQG equation is related to Onsager’s conjecture [Ons49]. As previously introduced, there are two types of energy that have physical significance for the SQG equation: the Hamiltonian and the kinetic energy. Isett and Vicol proved in [IV15_2d_Onsager] that $\theta\in L^{3}_{t}L^{3}_{x}$ implies the conservation of the Hamiltonian $\mathcal{H}$. For sufficiently regular solutions, the $L^{2}$-norm (the kinetic energy) is conserved as well. Several Besov-type regularity conditions are proposed for weak solutions to guarantee the conservation of the kinetic energy, see [Zho05_energy], [Cha06_energy], [WYY23_energy]. Dai [Dai17_energy] investigated the kinetic energy equality for viscosity solutions of the super-critical dissipative SQG equation under some regularity conditions.
@@
Weak-strong uniqueness for the SQG equation. Results of weak-strong uniqueness for the SQG equations with different fractional dissipation index $\alpha$ and different regularity conditions for initial data $\theta_{0}$ can be found in the following literature. Constantin and Wu [CW99_Weak_strong] proved the uniqueness of strong solution when $\alpha\in(1/2,1]$ with $\theta_{0}$ in $\dot{H}^{-1/2}$ or $L^{2}$ , which is analogous to the Ladyzhenskaya-Prodi-Serrin condition of the Navier-Stokes equation [KL57, Pro59, Ser62]. In the critical and supercritical cases $\alpha\in(0,1/2]$, there are several results for $L^{2}$ initial data. Dong and Chen in [DC06_weak_strong] proposed Ladyzhenskaya-Prodi-Serrin type condition $L^{r}_{t}L^{p}_{x}$ to $\nabla\theta$ and showed a weak-strong uniqueness property. In [DC12_weak_strong_uniqueness], the authors extended this regularity condition to Besov space $L^{r}_{t}B^{0}_{p,\infty}$. For the critical case $\alpha=1/2$, Marchand [Mar08_Weak_strong_critical] proved the weak-strong uniqueness in BMO-type space. Liu, Jia, and Dong [LJD12] proposed another BMO-type condition and proved the weak-strong uniqueness for all $\alpha\in(0,1).$ We will summarize examples of weak-strong uniqueness for subcritical, critical, and supercritical cases in Appendix A, and illustrate that these regularity classes can be taken as examples of the domain where we restrict the lower bound of large deviations. Furthermore, we point out that, to the best of our knowledge, weak-strong uniqueness results for the supercritical case $\alpha\in(0,1/2)$ with $\theta_{0}$ in $\dot{H}^{-1/2}$ have not been obtained. In this case, the dissipation term is not sufficient to prevent weak solutions from being distribution-valued. Therefore, it is challenging to propose a reasonable regularity condition for proving the uniqueness.
@@
Relationship to the study of convex integration. From the previous explanations, it can be inferred that obtaining matching upper and lower bounds for large deviations is related to the uniqueness of solutions for (1.1). Recently, the convex integration technique has led to various non-uniqueness results for fluid equations. Buckmaster, Shkoller, and Vicol [BSV19_Nonuniqueness_SQG] demonstrated the non-uniqueness of weak solutions for both inviscid and dissipative SQG equations by converting the equation into momentum form. Subsequently, different convex integration schemes have been employed to construct non-unique solutions for inviscid and dissipative SQG equations, as illustrated in [CKL21] and [IM21]. All of these works focus on the SQG equations without external forces, where the solutions satisfy $\Lambda^{-1}\theta\in C_{t}^{\sigma}C_{x}^{s}$ for some $\sigma$ and $s$. However, it remains an open question whether $\theta\in L^{2}_{t}H^{\alpha}_{x}$, which is the space considered in Theorem 1.4. For example, in [BSV19_Nonuniqueness_SQG], the solutions were constructed in $C_{t}^{\sigma}C_{x}^{s}$ with $1/2<s<4/5$ and $\sigma<s/(2-s).$ Therefore, these non-uniqueness results do not imply a violation of the full large deviations.
@@
In [HZZ23] and [HLZZ24], the authors discussed the non-uniqueness and non-Gaussianity for the stochastic SQG equations driven by irregular spatial noise and space-time white noise, respectively. We remark that, in [HLZZ24], infinitely many non-Gaussian ergodic stationary solutions of (1.3) (with $\delta=0$) in $B^{-1/2}_{p,1},p\geqslant 2,$ were constructed. We hope to connect the discussion of non-Gaussian stationary solutions with the uniqueness problem of deterministic SQG equations through the non-Gaussian large deviations in Section 8.
@@ 1.4. Structure of the paper
This paper is organized as follows. In Section 2, we introduce some preliminaries which will be used throughout the paper. Section 3 is dedicated to proving the existence and exponential tightness for stochastic generalized Leray solutions. Section 4 presents a variational characterization of the rate function. In Section 5, we establish the upper bound for large deviations. Section 6 provides a restricted lower bound for large deviations using the entropy method. In Section 7, we prove a relationship between the weak-strong uniqueness class and the energy equality for (1.1). Finally, in Section 8, we discuss an explicit representation of the quasi-potential. Additionally, Appendix A summarizes several examples of weak-strong uniqueness classes concerning different ranges of parameters $\alpha$ and $\beta$.
@@ 2.1. Notations and spaces
endowed with the norm $\|f\|_{W^{\sigma,p}([0,T];E)}^{p}:=\int_{0}^{T}\|f(t)\|_{E}^{p}\mathrm{\,d}t+\int_{0}^{T}\int_{0}^{T}\frac{\|f(t)-f(s)\|_{E}^{p}}{|t-s|^{1+\sigma p}}\mathrm{\,d}t\,\mathrm{d}s$. We denote the distribution space on $\mathbb{T}^{2}$ by $\mathscr{D}^{\prime}(\mathbb{T}^{2})$, which is the dual space of $C^{\infty}(\mathbb{T}^{2})$. Let $f\in\mathscr{D}^{\prime}(\mathbb{T}^{2})$ and let $g=(g(k))_{k\in\mathbb{Z}^{2}}$ be at most of polynomial growth. The Fourier transform of $f$ and the inverse Fourier transform of $g$ are defined as
$\mathcal{F}_{\mathbb{T}^{2}}f(k)$ $:=\hat{f}(k):=\int_{\mathbb{T}^{2}}e^{2\pi\mathrm{i}x\cdot k}f(x)\mathrm{\,d}x,\quad k\in\mathbb{Z}^{2},$
$\mathcal{F}_{\mathbb{T}^{2}}^{-1}g(x)$ $:=\sum_{k\in\mathbb{Z}^{2}}e^{2\pi\mathrm{i}x\cdot k}g(k),\quad x\in\mathbb{T}^{2}.$
Let $H:=\left\{f\in L^{2}\left(\mathbb{T}^{2}\right):\int_{\mathbb{T}^{2}}f\,\mathrm{d}x=0\right\}$ and let $\langle\cdot,\cdot\rangle$ denote the inner product on $H.$ Let $\mathbb{Z}_{+}^{2}:=\left\{\left(k_{1},k_{2}\right)\in\mathbb{Z}^{2}|k_{2}>0\right\}\cup\left\{\left(k_{1},0\right)\in\mathbb{Z}^{2}|k_{1}>0\right\}$ and $\mathbb{Z}_{-}^{2}:=\left\{\left(k_{1},k_{2}\right)\in\mathbb{Z}^{2}|-k\in\mathbb{Z}_{+}^{2}\right\},$ then
is an orthonormal eigenbasis of $\Lambda$ on $H$, which we denote by $\{e_{k}\}$. For every $s>0$, define $\|f\|_{H^{s}(\mathbb{T}^{2})}^{2}:=\sum_{k\in\mathbb{Z}^{2}}|2\pi k|^{2s}\big|\hat{f}(k)\big|^{2}$, and let $H^{s}(\mathbb{T}^{2})$ denote the space of all $f\in H$ for which $\|f\|_{H^{s}(\mathbb{T}^{2})}$ is finite. Define $H^{-s}(\mathbb{T}^{2})$ to be the dual of $H^{s}(\mathbb{T}^{2})$. Finally, let $H^{\infty}(\mathbb{T}^{2}):=\bigcap\limits_{s>0}H^{s}(\mathbb{T}^{2})$.
@@
$S_{j}f:=\mathcal{F}_{\mathbb{T}^{2}}^{-1}\left(\varphi\left(\cdot/2^{j}\right)\hat{f}\right)=\sum_{k=-\infty}^{j-1}\Delta_{k}f,\quad H_{j}f:=\left(Id-S_{j}\right)f=\sum_{k=j}^{\infty}\Delta_{k}f.$ (2.1)
For every $\phi\in H^{\infty}(\mathbb{T}^{2})$, the commutator between $\Lambda$ and $\phi$ is defined by $\Lambda,\phi]g:=\Lambda(\phi g)-\phi\Lambda g$, for any $g\in H^{\infty}(\mathbb{T}^{2}).$ The following lemma enables us to introduce the commutator estimate, which will contribute to the well-definedness of the nonlinear term.
@@
Lemma 2.1 can be proved by integration by parts, see for example, [Mar08_existence, Lemma 2.1].
@@
This commutator estimate is stated for $\mathbb{R}^{2}$ functions in [Mar08_existence, Lemma 2.2]. The same proof also works in the case of $\mathbb{T}^{2}$ functions, see [ZZ14, Lemma 4.2], [BSV19_Nonuniqueness_SQG, Lemma A.5].
@@
$\left\langle\nabla\cdot\left(H_{j}fR^{\perp}\left(H_{j}f\right)\right),\phi\right\rangle$= $-{}_{H^{-1/2}}\big\langle\frac{1}{2}R_{2}\left(H_{j}f\right),\left[\Lambda,\partial_{1}\phi\right]\Lambda^{-1}\left(H_{j}f\right)\big\rangle_{H^{1/2}}$ (2.3)
@@
2.2. The Regularization of Noise
[…]
A direct computation shows that
[…]
This indicates the blow-up speed of solutions of (1.3) and is therefore related to the scaling regimes we will encounter later on. This lead us to require that the scaling regime (1.7) holds for $(\varepsilon,\delta(\varepsilon))$ . Similarly, we require that the scaling regime $(\varepsilon,m(\varepsilon))$ for the Galerkin sequence (3.1) satisfies
In this context, we denote $\xi=\mathrm{d}W(t)/\mathrm{d}t$ and $\xi_{\delta(\varepsilon)}=\sqrt{Q_{\delta(\varepsilon)}}\mathrm{d}W(t)/\mathrm{d}t.$